OpenStudy (anonymous):

Write the point-slope equation of the line through (-2, 1) with a slope of 3.

OpenStudy (anonymous):

y-y1=m(x-x1) y-1=3[x-(-2)] y-1=3x+6

OpenStudy (anonymous):

A. y - 1 = 3(x + 2) B. y - 2 = 3(x + 1) C. y + 2 = 3(x - 1)

OpenStudy (anonymous):

its A i just simplified it more

OpenStudy (anonymous):

Thanks <3

OpenStudy (anonymous):

anytime